OpenStudy (alexa98):

Maria is playing a game where she is trying to draw a spade from a standard deck of cards. if she doesnt get a spade, she replaces the card, shuffles the deck, and tries again. if she draws a card 3 times and doesnt get a spade, she loses. what is the probability that Maria loses the game? round your answer to the nearest tenth of a percent.

satellite73 (satellite73):

not spade times not spade times not space

satellite73 (satellite73):

what is the probability she does not draw a spade?

satellite73 (satellite73):

if that is not clear, let me know

OpenStudy (rayep12):

27/64

OpenStudy (rayep12):

.4

OpenStudy (alexa98):

42.1%

satellite73 (satellite73):

you know where any of those numbers come from?

OpenStudy (alexa98):

wouldn't it be 39/52?

satellite73 (satellite73):

no there are four suits, with equal number of cards in each so the probability of not drawing a spade on any one try is \(\frac{3}{4}\)

satellite73 (satellite73):

of course you are also right, since \[\frac{39}{52}=\frac{3}{4}\] it is just that \(\frac{3}{4}\) is a heck of a lot easier to cube

OpenStudy (alexa98):

thank you!

satellite73 (satellite73):

yw

OpenStudy (rayep12):

42.1%